# Don’t Trust Generative Agents to Mimic Communication on Social Networks Unless You Benchmarked their Empirical Realism

### v1

@@ 1 Introduction
  • Formalizing social networks to $i)$ quantify and standardize their simulation and $ii)$ allow to benchmark their empirical realism.
@@ 2 Related Work
Although these applications show promise, they also raise significant methodological and ethical questions. The reliability and validity of using LLMs to represent human behavior and cognition is still subject to debate. Current research raises concerns about potential biases in training data that lead to misrepresentation of certain groups or points of view Abid et al. (2021); Hutchinson et al. (2020). Furthermore, the use of LLMs as replacements for human participants in social research raises ethical considerations about informed consent and the potential for misuse or misinterpretation of results. Without a deeper understanding of the model’s representations of ideologies, we risk oversimplifying complex human behaviors and social dynamics. Especially since these approaches ignore the LLMs lack of embodiment in the physical world Argyle et al. (2023). This disembodied nature means they lack the grounding in physical reality – expressed by cultural contexts, physical environments, and interpersonal relationships – that shapes human cognition, perception, and decision making Hussein (2012).
Although less complex than modeling communicative behavior, research on LLMs generating synthetic public opinion demonstrates whether base LLMs can accurately represent opinions in a sociocultural context. Although LLMs show an excessive ability to replicate US personas - closely approximating real world opinion distributions - they underperform when being tasked with replicating more complex non-English populations Argyle et al. (2023); Ma et al. (2024). Ma et al. demonstrate sensitivity to variable inclusion/exclusion. Similarly Tjuatja et al. (2024) report significant behavioral changes in LLM trained with RLHF due to question perturbations, attributing this behavior to the training scheme, stressing the need to adjust LLMs to case-specific data to improve grounding in reality.
@@ 3.1 Modeling Agents
Let $A$ be a set of agents, i.e., social media users, where $a_{t}^{i}$ is the state of the $i$-th agent at time $t$ [^3] The internal state of an agent can be represented by symbolic logic, machine learned functions, cognitive architecture or any other agent formalism that is capable of perceiving its environment and acting upon that environment.
$M$ is a set of messages, where $m_{t}^{(i,j)}$ is a message from agent $i$ to agent $j$ at time $t$. Consequently, $m_{t}^{(i,\cdot)}$ indicates a message that agent $i$ transmits to all other agents and $M_{t}^{(\cdot,i)}$ indicates all messages sent to agent $i$ [^4] Note, that messages can be of any modality or type (including images, videos, likes, shares,…) Accordingly, $M_{t}^{i}$ indicates all messages that agent $i$ sends and receives at time $t$.
@@ 4 Simulating Online Social Networks
Behavior $b$ Agent $A$ Task Type Input $\widetilde{M}_{t}^{(\cdot,i)}$ Output $m_{t+1}^{(i,\cdot)}$ Loss $L_{b}$ Metric(s) $q$
@@ 6 Fitting Agents to User Data from $\mathbb{X}$
The complete technical pipeline for this study is publicly available on GitHub at https://github.com/cl-trier/TWON-Agents. This repository includes the source code for all the data processing, training, and evaluation steps described in this paper. The raw and processed datasets used in this study will be made available after acceptance.
@@ 7.2 Evaluation of Replying Likelihood Behavior
English Ignored 0.978 0.978 0.978 500
Replied 0.978 0.978 0.978 500
Weighted Avg 0.978 0.978 0.978 1000

### v2

@@ 1 Introduction
  • Formalizing social networks to $i)$ quantify and standardize their simulation and ii)ii) allow to benchmark their empirical realism.
@@ 2 Related Work
Although these applications show promise, they also raise significant methodological and ethical questions. The reliability and validity of using LLMs to represent human behavior and cognition is still subject to debate Larooij and Törnberg (2025). Current research raises concerns about potential biases in training data that lead to misrepresentation of certain groups or points of view Abid et al. (2021); Hutchinson et al. (2020). Furthermore, the use of LLMs as replacements for human participants in social research raises ethical considerations about informed consent and the potential for misuse or misinterpretation of results. Without a deeper understanding of the model’s representations of ideologies, we risk oversimplifying complex human behaviors and social dynamics. Especially since these approaches ignore the LLMs lack of embodiment in the physical world Argyle et al. (2023). This disembodied nature means they lack the grounding in physical reality – expressed by cultural contexts, physical environments, and interpersonal relationships – that shapes human cognition, perception, and decision making Hussein (2012).
Although less complex than modeling communicative behavior, research on LLMs generating synthetic public opinion demonstrates whether base LLMs can accurately represent opinions in a sociocultural context. Although LLMs show an excessive ability to replicate US personas - closely approximating real world opinion distributions - they underperform when being tasked with replicating more complex non-English populations Argyle et al. (2023); Ma et al. (2024). Ma et al. demonstrate sensitivity to variable inclusion/exclusion. Similarly Tjuatja et al. report significant behavioral changes in LLM trained with RLHF due to question perturbations, attributing this behavior to the training scheme, stressing the need to adjust LLMs to case-specific data to improve grounding in reality.
@@ 3.1 Modeling Agents
Let $A$ be a set of agents, i.e., social media users, where $a_{t}^{i}$ is the state of the $i$-th agent at time tt. [^3] The internal state of an agent can be represented by symbolic logic, machine learned functions, cognitive architecture or any other agent formalism that is capable of perceiving its environment and acting upon that environment.
$M$ is a set of messages, where $m_{t}^{(i,j)}$ is a message from agent $i$ to agent $j$ at time $t$. Consequently, $m_{t}^{(i,\cdot)}$ indicates a message that agent $i$ transmits to all other agents and $M_{t}^{(\cdot,i)}$ indicates all messages sent to agent ii. [^4] Note, that messages can be of any modality or type (including images, videos, likes, shares,…) Accordingly, $M_{t}^{i}$ indicates all messages that agent $i$ sends and receives at time $t$.
@@ 4 Simulating Online Social Networks
Behavior $b$ Agent AA Task Type Input $\widetilde{M}_{t}^{(\cdot,i)}$ Output $m_{t+1}^{(i,\cdot)}$ Loss $L_{b}$ Metric(s) $q$
@@ 6 Fitting Agents to User Data from $\mathbb{X}$
The complete technical pipeline for this study is publicly available on GitHub at https://anonymous.4open.science/r/TWON-Agents-D0E2. This repository includes the source code for all the data processing, training, and evaluation steps described in this paper. The raw and processed datasets used in this study will be made available after acceptance.
@@ 7.2 Evaluation of Replying Likelihood Behavior
English Ignored 0.973 0.972 0.972 500
Replied 0.972 0.974 0.973 500
Weighted Avg 0.973 0.973 0.972 1000
@@ 8.2 Future Research
Acknowledgments
This study was conducted with a financial contribution from the EU’s Horizon Europe Framework (HORIZON-CL2-2022-DEMOCRACY-01-07) under grant agreement number 101095095.